Calculate the number of H atoms present in a 5.00 g sample of NH3 (g)?
Answer #1

solution: The formula mass/molar mass of Nitza =14+1X3 = 17 Imol * No. of moles of NH3 = 59 - 179/mol = 0.294 moles since, o